# DiagramQG: A Dataset for Generating Concept-Focused Questions from Diagrams

### v1

@@ Abstract
Visual Question Generation (VQG) has gained significant attention due to its potential in educational applications. However, VQG researches mainly focus on natural images, neglecting diagrams in educational materials used to assess students’ conceptual understanding. To address this gap, we introduce DiagramQG, a dataset containing 8,372 diagrams and 19,475 questions across various subjects. DiagramQG introduces concept and target text constraints, guiding the model to generate concept-focused questions for educational purposes. Meanwhile, we present the Hierarchical Knowledge Integration framework for Diagram Question Generation (HKI-DQG) as a strong baseline. This framework obtains multi-scale patches of diagrams and acquires knowledge using a visual language model with frozen parameters. It then integrates knowledge, text constraints and patches to generate concept-focused questions. We evaluate the performance of existing VQG models, open-source and closed-source vision-language models, and HKI-DQG on the DiagramQG dataset. Our HKI-DQG outperform existing methods, demonstrating that it serves as a strong baseline. Furthermore, we apply HKI-DQG to two other VQG datasets of natural images, namely VQG-COCO and K-VQG, achieving state-of-the-art performance. The dataset and code are available at https://dxzxy12138.github.io/diagramqg-home.
@@ 1 Introduction
Visual Question Generation (VQG), a crucial research area in computer vision and natural language processing, aims to automatically generate meaningful questions from input images. The evolution of VQG approaches has witnessed a significant shift from traditional rule-based and template-based methodologies to sophisticated neural network architectures, encompassing both unconstrained [27] and constrained [14, 11, 35, 34] paradigms. While early VQG research focused primarily on generating questions using only image inputs [27], contemporary research has expanded to incorporate various guiding information to enhance question generation specificity, including answers [14, 16], answer types [11, 36], image region [35], question types [7], and knowledge triple [34], thereby significantly improving the quality and relevance of generated questions. We demonstrate various existing text constraints as inputs for visual question generation, as shown in (a) of Figure 1.
@@
However, current VQG research focuses on natural images, neglecting diagrams commonly popular in educational materials. This fails to meet educational requirements. Diagrams play a vital role in education by presenting information in a structured engaging manner, enabling students to understand complex concepts more effectively [4]. Moreover, these concept-focused questions based on diagrams assess students’ ability to explain, analyze, and apply the information presented in diagrams using concepts [12]. Therefore, generating questions based on diagrams is crucial for assessing students’ ability to apply the learned concepts in a practical context.
@@
To bridge the gap between current research and educational requirements, we have constructed DiagramQG, the first dataset containing 8,372 diagrams and 19,475 questions, covering 4 subjects, 15 courses, and 169 concepts, by integrating existing diagram question-answering datasets [10, 40, 49, 23] with secondary school curriculum materials. These questions provide a more comprehensive evaluation of students’ understanding of the concepts across different subjects. Meanwhile, the DiagramQG dataset introduces concept and target text constraints, guiding the model to generate questions that meet educational requirements and include the target text, as shown in (b) of Figure 1. The target text serves as guidance to control question diversity, as each diagram can generate multiple concept-focused questions. Related research shows that answer-based constraints tend to generate context-independent questions [18]. Question-type constraints often deviate from the intended assessment objectives [34]. Moreover, knowledge triple-based constraints frequently result in mere triple expansions rather than meaningful questions [25]. This demonstrates the importance of our proposed text constraint for diagram question generation.
@@
Compared to natural images, diagrams require a significant amount of subject-specific background knowledge for proper understanding [41], and they typically contain a substantially higher number of objects [9]. Thus, tackling the Diagram Question Generation (DQG) task requires a framework that can incorporate subject-specific knowledge, comprehend complex object relationships within diagrams, and generate meaningful questions accordingly. We propose a novel Hierarchical Knowledge Integration framework for Diagram Question Generation (HKI-DQG). The framework uses CLIP to identify relevant diagram patches of various scales based on text constraints. Advanced vision-language models like BLIP and Qwen2-VL extract knowledge from these multi-scale patches. T5 then filters the extracted knowledge to ensure relevance to the text constraints. Finally, the framework combines the text constraints, selected knowledge, and multi-scale diagram patches to generate questions. Notably, this approach doesn’t require training the vision-language model, which improves scalability and reduces costs. We evaluate the performance of current visual question generation models, open-source vision-language models, closed-source vision-language models, and HKI-DQG on our constructed DiagramQG dataset. Our contributions are summarized as follows:
@@
  • We introduce DiagramQG, the first question generation dataset specifically designed for diagrams, and develop a novel text constraint that guides the model in generating concept-focused questions.
@@
  • We propose an adaptable framework for DQG based on different frozen-parameter vision-language models, addressing DQG tasks at a lower cost.
@@
  • We evaluate the performance of current mainstream VQG models, various open-source and closed-source vision-language models, and our proposed framework on the constructed DiagramQG dataset.
@@ 2.1 Visual Question Generation
VQG task has progressed from rule-based and template-based methods to advanced neural network architectures, including both unconstrained and constrained approaches. The unconstrained approach, while simple to implement, often produces generic questions with insufficient image specificity, restricting the development of constrained VQG methods [1, 27]. Researchers have explored various text constraints to improve question generation quality, including answer-guided methods [11, 43, 46, 20], question types [7], answer categories [38], and knowledge-enhanced approaches [44, 2]. Recent works have proposed cross-topic VQG models [19] and dual contrastive learning approaches [25] to address unseen topics and improve visual-textual alignment. However, current methods face challenges in balancing question diversity and effective utilization of visual information, emphasizing the need for innovative text constraints and methods that ensure constraint validity, question naturalness, and diversity while maximizing the use of visual information.
@@ 2.2 Diagram Question Answering
Diagram Question Answering (DQA) requires higher reasoning ability and background knowledge compared to Visual Question Answering (VQA). Previous research has focused on enhancing diagram understanding through various pre-training approaches [8, 24, 45]. Additionally, studies have investigated the use of prompting methods with Large Language Models (LLMs) to harness their extensive prior knowledge for DQA [23, 50, 47, 41]. Aligning the information in the diagram with the required background knowledge is crucial for understanding the diagram and achieving effective DQA performance.
@@ 3 The DiagramQG Dataset
The DiagramQG dataset is a comprehensive collection of questions covering 4 subjects, 15 courses, and 169 concepts, consisting of 19,475 unique questions and 8,372 diagrams. These components form 44,472 combinations of tuples (target & concept text constraint, diagram, question). We provide examples from four different subjects in the DiagramQG dataset, as shown in Figure 2. The diagram list annotates objects appearing in the diagram, which will be used in later evaluation metrics.
@@ 3.1 Data Collection
The data collection process is divided into four distinct phases. The data collection process begins by gathering content from well-established datasets [10, 40, 49, 23], supplemented by Google Image Search. This initial phase yields a substantial raw dataset containing over 20,000 diagrams and 40,000 questions spanning multiple subjects.
Subsequently, the collected materials are organized into four primary subjects and further categorized into 15 specific courses. This structuring process involves mapping questions to their corresponding concepts, resulting in the identification of 169 distinct concepts.
The third phase involves comprehensive annotation work conducted by trained crowd workers, who annotate target & concept text constraints for each diagram-question pair. This annotation effort produces more than 70,000 unique combinations, as some questions are associated with different target & concept text constraints, such as the same concept but different targets. Additionally, they annotate diagram elements and texts to evaluate models’ diagram comprehension capabilities.
The final quality assurance phase employs a separate group of crowd workers to evaluate each combination using a 0-100 scale with 5-point intervals. After filtering out combinations scoring below 60 points, the curated dataset comprises 8,372 diagrams, 19,475 questions, and 44,472 complete combinations of diagrams, questions, and their associated target and concept text constraints.
@@ 3.2 Data Analyse
Num. of Q Num. of I Object on I Images Text Constraint Knowledge types
@@
FVQA 5,826 2k 2.9 natural answer common-sense
VQG-COCO 25,000 5k 3.3 natural image, caption common-sense
K-VQG 16,098 13K 2.7 natural knowledge triple common-sense
DiagramQG 19,475 8,372 11.2 diagram target, concept subject knowledge
Table 1: Comparing core characteristics of major VQG Datasets and DiagramQG
@@
The DiagramQG dataset encompasses a broad spectrum of academic domains, systematically categorized into four primary subjects: Natural Science (blue), Earth Science (yellow), Applied Science (green), and Social Science (orange), as shown in Figure 3. The dataset is structured hierarchically, with samples first classified by subject, then further divided into specific courses (e.g., Biology), and ultimately categorized into detailed concepts (e.g., Ecological interactions). Figure 3 presents a treemap visualization illustrating the dataset’s breadth and depth, encompassing 4 subjects, 15 courses, and 169 concepts across several academic domains.
The DiagramQG dataset is distinguished by its diverse range of scientific questions, which exhibit significant variation in both question formulation and complexity. As depicted in Figure 4, an analysis of the first four words in question texts reveals a wide array of question types and structures. The diversity is further exemplified by the variation in question length, ranging from 4 to 50 words, with an average length of 9.52 words, underscoring the comprehensive nature of the dataset.
@@
Table 1 presents a comprehensive comparison between the DiagramQG dataset and existing visual question generation datasets. The comparative analysis reveals that DiagramQG distinguishes itself by focusing on specialized concepts across multiple subjects, in contrast to other datasets that primarily emphasize general diagram descriptions and common sense reasoning. Moreover, DiagramQG stands out as the first large-scale multimodal scientific question generation dataset that incorporates non-answer words or phrases and concepts across different subjects as text constraints. A notable aspect of our analysis involves the examination of object density per diagram across the five datasets. This is particularly significant considering that diagrams inherently contain more complex and information-rich content compared to natural images.
@@ 3.2.3 Challenges in DiagramQG dataset
Based on our comparative analysis, DiagramQG presents three unique challenges that differentiate it from existing visual question generation datasets:
  • Domain-specific knowledge requirement: Unlike other datasets that rely on general common sense, DiagramQG encourages models to possess and apply different concepts across various subjects to generate meaningful and practical questions.
  • Long-tail distribution phenomenon: The long-tail distribution in DiagramQG, where concept coverage ranges from abundant to limited, challenges models to generalize and optimize performance across both well-represented and underrepresented concepts.
  • High object information density: The high density of object information in diagrams complicates content interpretation and risks overlooking critical details, demanding models to possess capabilities in capturing and processing complex visual information.
@@ 4.1 Problem Definition
The Diagram Question Generation (DQG) task aims to generate a reasonable question $q$ given a diagram $d$, a target text $t$, and a concept text $c$. The generated question should contain the target text $t$ and effectively assess students’ understanding of the specified concept $c$. This task can be formulated as a conditional generation problem, represented as $p(q|d,t,c)$, where a multimodal model maps visual and textual information into a joint embedding space before decoding questions that satisfy both the text constraint and the conceptual assessment requirement.
@@ 4.2.1 HierKnowExtract
For textual components, the T5 Encoder [33] is employed to process both the target text $t$ and concept text $c$ independently. We introduce a learnable linear transformation $W_{h}$ to ensure compatibility between the CLIP Image Encoder [32] and T5 Encoder [33] feature spaces. This facilitates unified similarity computation and subsequent operations in the KnowFusionQG phase. The process culminates in the selection of the most semantically relevant patch from each hierarchical layer to form the set $P_{s}$, as shown in Equation 2.
@@
Selected patches $P_{s}$, target text $t$, and concept text $c$ are fed into a large-scale vision-language model (VLM) like Qwen-VL or BLIP to obtain diverse patch-specific knowledge. The resulting knowledge set $K_{s}$ is generated according to Equation 3.
@@ 4.2.2 KnowSelect
The KnowSelect stage selects the $n$ most relevant knowledge sentences $K_{(t,c)}$ from the extensive knowledge set $K_{s}$ based on target text $t$ and concept text $c$. The process begins with encoding the knowledge set $K_{s}$ and the text constraints using the T5 Encoder [33], as formulated in Equation 4.
@@
where $d_{k}$ denotes the dimensionality of the hidden representations. The top-m operator selects the m most semantically relevant knowledge sentences based on attention scores, creating a knowledge set $K_{(t,c)}$ that best matches the text constraints.
@@ 4.2.3 KnowFusionQG
The KnowFusionQG stage integrates the selected diagram patches $P_{s}$, target text $t$, concept text $c$, and knowledge set $K_{(t,c)}$ to generate the concept-focused question through a multimodal fusion mechanism. The process begins with encoding inputs into language and vision representations, as shown in Equation 6.
@@
Finally, the fused output $H_{fuse}$ is fed into the T5 decoder [33] to predict the target question $\hat{q}$. This integration of visual and textual information through cross-modal attention and gated fusion enables the model to generate questions that are both contextually relevant and conceptually focused.

### v3

# DiagramQG: Concept-Focused Diagram Question Generation via Hierarchical Knowledge Integration

@@ Abstract
Visual Question Generation (VQG) has gained significant attention due to its potential in educational applications. However, VQG research mainly focuses on natural images, largely neglecting diagrams in educational materials used to assess students’ conceptual understanding. To address this gap, we construct DiagramQG, a dataset containing 8,372 diagrams and 19,475 questions across various subjects. DiagramQG introduces concept and target text constraints, guiding the model to generate concept-focused questions for educational purposes. Meanwhile, we present the Hierarchical Knowledge Integration framework for Diagram Question Generation (HKI-DQG) as a strong baseline. This framework obtains multi-scale patches of diagrams and acquires knowledge using a visual language model with frozen parameters. It then integrates knowledge, text constraints, and patches to generate concept-focused questions. We evaluate the performance of existing VQG models, open-source and closed-source vision-language models, and HKI-DQG on the DiagramQG dataset. Our novel HKI-DQG consistently outperforms existing methods, demonstrating that it serves as a strong baseline. Furthermore, we apply HKI-DQG to four other VQG datasets of natural images, namely VQG-COCO, K-VQG, OK-VQA, and A-OKVQA, achieving state-of-the-art performance.
@@ 1 Introduction
Visual Question Generation (VQG) is a crucial research area in computer vision and natural language processing. It aims to automatically generate meaningful questions from input images. The evolution of VQG approaches has witnessed a significant shift from traditional rule-based and template-based methodologies to sophisticated neural network architectures, encompassing both unconstrained [26] and constrained [13, 11, 34, 33] paradigms. While early VQG research focused primarily on generating questions using only image inputs [26], contemporary research has expanded to incorporate various guiding information to enhance question generation specificity, including answers [13, 15], answer types [11, 35], image region [34], question types [7], and knowledge triple [33], thereby significantly improving the quality and relevance of generated questions. We demonstrate various existing text constraints as inputs for visual question generation, as shown in (a) of Figure 1.
@@
However, current VQG research focuses on natural images, neglecting diagrams commonly popular in educational materials. This fails to meet educational requirements for assessment. Diagrams play a vital role in education by presenting information in a structured engaging manner, enabling students to understand complex concepts more effectively [4]. Moreover, these concept-focused questions based on diagrams assess students’ ability to explain, analyze, and apply the information presented in diagrams using concepts [12] on diverse subjects. Therefore, generating questions based on diagrams is crucial for assessing students’ ability to apply the learned concepts in a practical context.
@@
To bridge the gap between current research and educational requirements, we have constructed DiagramQG, the first dataset containing 8,372 diagrams and 19,475 questions, covering 4 subjects, 15 courses, and 169 concepts, by integrating existing diagram question-answering datasets [10, 39, 48, 22] with curriculum materials we collected. These questions provide a more comprehensive evaluation of student’s understanding of the concepts across different subjects. Meanwhile, the DiagramQG dataset introduces concept and target text constraints, guiding the model to generate questions that assess students’ understanding of concepts and include the target text, as shown in (b) of Figure 1. The target text serves as guidance to control question diversity, as each diagram can generate multiple concept-focused questions. Related research shows that answer-based constraints tend to generate context-independent questions [17]. Question-type constraints often deviate from the intended assessment objectives [33]. Moreover, knowledge triple-based constraints frequently result in mere triple expansions rather than meaningful questions [24]. This highlights the need for novel text constraints to generate educationally meaningful questions.
@@
Compared to natural images, diagrams require a significant amount of subject-specific background knowledge for understanding [40], and they typically contain a higher number of objects [9]. Thus, tackling the Diagram Question Generation (DQG) task requires a framework that can incorporate subject-specific knowledge, comprehend object relationships within diagrams, and generate meaningful questions accordingly. We propose a novel Hierarchical Knowledge Integration framework for Diagram Question Generation (HKI-DQG). The framework leverages CLIP to identify relevant multi-scale diagram patches, from which vision-language models like BLIP and Qwen2-VL extract knowledge. After T5 filters the extracted knowledge to ensure relevance to text constraints, the framework integrates these filtered insights with text constraints and multi-scale diagram patches to generate questions. Notably, this approach doesn’t require training the vision-language model, which improves scalability and reduces costs. We evaluate HKI-DQG against current VQG and vision-language models on our DiagramQG dataset. Additionally, we conduct experiments with HKI-DQG on four natural images of VQG datasets, validating HKI-DQG’s generalizability and effectiveness. Our contributions are summarized as follows:
@@
  • We introduce the comprehensive DiagramQG dataset, the first question generation dataset specifically designed for diagrams, and develop a novel text constraint that guides the model in generating concept-focused questions.
@@
  • We propose an adaptable framework based on different frozen-parameter vision-language models, addressing DQG tasks at a lower cost, and conduct experiments on both DiagramQG and four natural image VQG datasets.
@@
  • We thoroughly evaluate the performance of current mainstream VQG models, various open-source and closed-source vision-language models, and our proposed framework on the newly constructed DiagramQG dataset.
@@ 2.1 Visual Question Generation
VQG task has progressed from rule-based and template-based methods to advanced neural network architectures, including both unconstrained and constrained approaches. The unconstrained approach often produces generic questions with insufficient image specificity, restricting the development of unconstrained VQG methods [1, 26]. Researchers have explored various text constraints to improve question generation quality, including answer-guided methods [11, 42, 45, 19], question types [7], answer categories [37], and knowledge-enhanced approaches [43, 2]. Recent works have proposed cross-topic VQG models [18] and dual contrastive learning approaches [24] to address unseen topics and improve visual-textual alignment. However, current methods face challenges in balancing question diversity and effective utilization of visual information, emphasizing the urgent need for innovative text constraints and robust methods that ensure constraint validity, question naturalness, and diversity in the use of visual information.
@@ 2.2 Diagram Question Answering
Diagram Question Answering (DQA) requires higher reasoning ability and background knowledge compared to Visual Question Answering (VQA). Previous research has focused on enhancing diagram understanding through pre-training approaches [8, 23, 44]. Additionally, studies investigate the use of prompting methods with Large Language Models to harness their extensive prior knowledge for DQA [22, 49, 46, 40]. Aligning the information in the diagram with background knowledge is crucial for understanding the diagram and achieving effective DQA performance.
@@ 3 The DiagramQG Dataset
The DiagramQG is a dataset of questions covering 4 subjects, 15 courses, and 169 concepts, consisting of 19,475 unique questions and 8,372 diagrams. These components form 44,472 combinations of tuples (target & concept text constraint, diagram, question). We provide examples from four subjects in the DiagramQG dataset, as shown in Figure 2. The diagram list annotates objects appearing in the diagram, which will be used in later evaluation metrics.
@@ 3.1 Data Collection
The data collection process can be divided into four distinct phases. This process begins by gathering content from well-established datasets [10, 39, 48, 22], supplemented by searching for relevant images on Hugging Face, GitHub, and Roboflow, with all data following CC BY-SA 4.0, CC BY-NC-SA, and MIT licenses. This initial phase yields a substantial raw dataset containing over 25,000 diagrams and 60,000 questions spanning multiple subjects.
Subsequently, the collected diagram and questions are organized into four primary subjects and further categorized into 15 specific courses. This structuring process involves mapping the questions to their corresponding concepts, resulting in the identification of 169 distinct concepts.
Next, experienced teachers from different academic subjects with teaching experience annotate target text constraints for each diagram-question pair under their specialized concepts. This detailed annotation effort produces more than 70,000 unique combinations, as some questions are associated with different target & concept text constraints, such as the same concept but different targets.
Finally, another group of subject teachers screens all current combinations based on their educational utility, using a 100-point scale with a passing threshold of 60 points. The resulting DiagramQG dataset contains 8,372 diagrams, 19,475 questions, and 44,472 complete combinations, sourced from ScienceQA (34.1%), TQA (24.3%), CSDQA (10.8%), ADE (9.2%), and our collection (21.6%).
@@ 3.2 Data Analyse
Num. of Q. Num. of I. Object on I. Images Text Constraint Science subjects
@@
FVQA 5,826 2k 2.9 natural answer N/A
VQG-COCO 25,000 5k 3.3 natural image, caption N/A
K-VQG 16,098 13K 2.7 natural knowledge triple N/A
ScienceQA 21,208 10,332 4.3 diagram answer natural, social
CSDQA 3,494 1,294 5.2 diagram answer computer science
TQA 26,260 3,455 6.8 diagram answer natural, earth
ADE 3.945 3.945 7.2 diagram answer natural
DiagramQG 19,475 8,372 11.2 diagram target, concept natural, earth, application, social
Table 1: Comparing core characteristics of major VQG datasets, DQA datasets and DiagramQG, where Q. and I. mean question and image.
@@
The DiagramQG dataset encompasses a broad spectrum of educational domains, categorized into four primary subjects: Natural Science (blue), Earth Science (yellow), Applied Science (green), and Social Science (orange), as shown in Figure 3. The dataset is structured hierarchically, with samples first classified by subject, then further divided into specific courses (e.g., Biology), and ultimately categorized into concepts (e.g., Ecological interactions). Figure 3 presents a treemap visualization illustrating the dataset’s breadth and depth, encompassing 4 subjects, 15 courses, and 169 concepts across several academic domains.
The DiagramQG dataset is distinguished by its diverse range of scientific questions, which exhibit significant variation in both question formulation and complexity. As clearly depicted in Figure 4, an extensive analysis of the first four words in question texts reveals a wide array of question types and structures. The remarkable diversity is further exemplified by the variation in question length, ranging from 4 to 50 words, with an average length of 9.52 words, underscoring the comprehensive nature of the dataset.
@@
Table 1 presents a comprehensive comparison between the DiagramQG dataset and existing VQG and DQA datasets. The comparative analysis reveals that DiagramQG demonstrates greater educational utility compared to VQG datasets, which primarily focus on general image description and common-sense reasoning tasks, and surpasses existing DQA datasets in subject matter coverage. Moreover, DiagramQG stands out as the first large-scale multimodal scientific question generation dataset that incorporates non-answer words or phrases and concepts across different subjects as text constraints. A notable aspect of our analysis involves the examination of object density per diagram across the nine datasets. This is particularly significant considering that diagrams inherently contain more complex and information-rich content compared to natural images.
@@ 3.2.3 Challenges in DiagramQG dataset
Based on our comprehensive comparative analysis, DiagramQG presents three unique challenges that differentiate it from existing visual question generation datasets:
  • Domain-specific knowledge requirement: Unlike other existing VQG datasets that rely on general common sense, DiagramQG consistently encourages models to possess and apply different concepts across various subjects to generate meaningful and practical questions.
  • Long-tail distribution phenomenon: The inherent complex long-tail distribution in DiagramQG, where concept coverage ranges from abundant to limited, challenges models to generalize and optimize performance across both well-represented and underrepresented concepts.
  • High object information density: The significantly high density of object information in the diagrams complicates content interpretation and risks overlooking critical details, demanding models to possess capabilities in capturing and processing complex visual information.
@@ 4.1 Problem Definition
The Diagram Question Generation (DQG) task aims to generate a concept-focused question $q$ given a diagram $d$, a target text $t$, and a concept text $c$. The generated question should contain the target text $t$ and effectively assess students’ understanding of the specified concept $c$. This task can be formulated as a conditional generation problem, represented as $p(q|d,t,c)$, where a multimodal model maps visual and textual information into a joint embedding space before decoding questions that satisfy both the text constraint and the conceptual assessment requirement.
@@ 4.2.1 HierKnowExtract
For textual components, the T5 Encoder [32] is employed to process both the target text $t$ and concept text $c$ independently. We introduce a learnable linear transformation $W_{h}$ to ensure compatibility between the CLIP Image Encoder [31] and T5 Encoder [32] feature spaces. This effectively facilitates unified similarity computation and subsequent operations in the KnowFusionQG phase. The process ultimately culminates in the selection of the most semantically relevant patch from each hierarchical layer to form the set $P_{s}$, as shown in Equation 2.
@@
Selected patches $P_{s}$, target text $t$, and concept text $c$ are carefully fed into a large-scale vision-language model (VLM) like Qwen2-VL or BLIP to obtain diverse patch-specific knowledge. The resulting knowledge set $K_{s}$ is systematically generated according to Equation 3.
@@ 4.2.2 KnowSelect
The KnowSelect stage selects the $m$ most relevant knowledge sentences $K_{(t,c)}$ from the extensive knowledge set $K_{s}$ based on target text $t$ and concept text $c$. The process begins with encoding the knowledge set $K_{s}$ and the text constraints using the T5 Encoder [32], as formulated in Equation 4.
@@
where $d_{k}$ denotes the dimensionality of the hidden representations. The top-m() operator selects the $m$ most semantically relevant knowledge sentences based on attention scores, creating a knowledge set $K_{(t,c)}$ that best matches the text constraints.
@@ 4.2.3 KnowFusionQG
The KnowFusionQG stage integrates the selected diagram patches $P_{s}$, target text $t$, concept text $c$, and knowledge set $K_{(t,c)}$ to generate the concept-focused question through a multimodal fusion mechanism. This process begins with encoding the visual and textual inputs into language and vision representations, as shown in Equation 6.
@@
Finally, the fused output $H_{fuse}$ is fed into the T5 decoder [32] to predict the target question $\hat{q}$. This integration of visual and text information through cross-modal attention and gated fusion enables the model to generate questions that are contextually relevant and conceptually focused.